OpenStudy (anonymous):

Let f(x)=x+4 and g(x)=x^2+8x –3. Find each of the following functions: a:(f+g)(x) b:(f –g)(x) c:(fg)(x) d:f∘g

OpenStudy (anonymous):

(f+g) just add the two functions, you get x^2 + 9x + 1

OpenStudy (anonymous):

(f-g), subtract , you get x+4 - (x^2 + 8x - 3) = x + 4 -x^2 - 8x + 3 = -x^2 -7x+7

OpenStudy (anonymous):

(fg)...multiply the two functions, you get (x+4)(x^2 + 8x - 3)...now just multiply the two

OpenStudy (anonymous):

(f/g)..divide the two functions, you get (x+4)/(x^2+8x-3)

OpenStudy (anonymous):

the last part is f*g